A brand of lightbulb has an average life of 600 hr, with a standard deviation of 40 hr. The length of life of the bulb can be closely approximated by a normal curve. A warehouse manager buys and installs​ 10,000 such bulbs. Find the total number that can be expected to last less than

716 hr

The total number of bulbs that can be expected to last less than

716 hr is

​(Round to the nearest whole number as​ needed.)
Answer #1

μ = 600

σ = 40

P( X < 716)

Z = X - μ / (σ / √n)

Z = 716 - 600 /( 40)

Z = 2.9

P( z < 2.9)

= 0.5 + 0.4981

= 0.9981

Number of bulbs that last less than 716 hr

= 0.9981 x 10000

= 9981